# for Verifiable Chain-of-Thought Financial Reasoning

### v2

@@ 4.3 Final Answer Correctness
Beyond step-level reasoning alignment, we also assess the correctness of the final predicted outcome. Let $s^{*}_{m}$ and $\hat{s}_{n}$ denote the last steps of the gold and predicted solutions, respectively. We define the Final Answer Correctness (FAC) metric as:
@@
using the same tolerance $\epsilon=0.05$ as before. FAC measures whether the model’s final computation aligns with the correct end result, complementing the DTW-based metric that evaluates reasoning faithfulness throughout the entire solution sequence.
Finally, ChainEval equals $\mathrm{DTWNormGate}$. It captures both local semantic and numeric agreement as well as global sequence-level coherence in the reasoning process. We verify that this metric best reflects the true reasoning quality by comparing it with human evaluations, as detailed in § 5.2.
@@ 5.1 Evaluated Models
We evaluate a total of 26 LLMs, grouped into four categories according to their capability and relevance to financial reasoning. (1) Frontier proprietary models represent the current upper bound of performance, including GPT-{5, 4.1, 5-mini, 4.1-mini} (OpenAI, 2025a, b), Claude Sonnet {4.5, 4, 3.7} (Anthropic, 2025b, c, a), Gemini 2.5 {Pro, Flash} (Comanici et al., 2025), DeepSeek {V3.2, V3.1, R1} (Liu et al., 2024; Guo et al., 2025), and Grok 4 {Heavy, Fast} (xAI, 2025). (2) Finance-specific models include Fino1 (Qian et al., 2025), FinR1 (Liu et al., 2025), DianJin-R1 (Zhu et al., 2025) and WiroAI Finance-{LLaMA, Qwen} (Abdullah Bezir, 2025), which are fine-tuned on financial corpora and reasoning templates to enhance domain accuracy. (3) Math-enhanced models include WizardMath (Luo et al., 2023), MetaMath (Yu et al., 2023), Mathstral (Mistral, 2024), and Qwen2.5-Math (Yang et al., 2024), which are trained on mathematical and symbolic datasets to improve quantitative reasoning. (4) General-purpose open models include LLaMA 3.1 (Grattafiori et al., 2024) and Qwen {2.5, 3} (Qwen, 2024, 2025), serving as strong domain-agnostic backbones for many specialized variants. This four-way taxonomy enables a structured comparison across reasoning paradigms, highlighting how model scale, domain specialization, and mathematical supervision jointly influence symbolic financial reasoning. Detailed configurations and model sources are provided in Appendix E.
@@ 5.2 ChainEval Validation
Before conducting large-scale experiments, we validated the proposed ChainEval through a controlled expert evaluation. We randomly sampled 20 instantiated questions from the dataset and generated answers using five models of different capacities and training paradigms, namely GPT-5, GPT-4.1 mini, MetaMath, Fino1, and LLaMA 3.1, to ensure a clear range of reasoning quality. This process produced 100 model-generated responses, which were then randomized and anonymized for human assessment. Financial experts independently evaluated each response with respect to the corresponding question and gold-standard reasoning trace. Each output was rated along two dimensions, Reasoning Process Quality and Final Answer Accuracy, on a five-point scale as described in Appendix F. We observed a strong association between these dimensions, with Spearman’s $\rho$ exceeding 0.94, confirming that coherent reasoning typically leads to accurate final answers. Consequently, subsequent analyses focus on Reasoning Process Quality as the primary evaluation dimension. We further computed several variants of ChainEval and compared their correlations with expert process scores against reference-based metrics such as ROUGE-2 and ROUGE-L (Lin, 2004). Among all variants, the DTWNormGate formulation showed the highest correlation with human judgments, capturing both semantic and numeric consistency across reasoning steps. Therefore, DTWNormGate is adopted as the principal evaluation metric in this study, with detailed ablation results reported in Appendix G.
@@ 5.3 Experimental Setup
We instantiate the benchmark by sampling 10 instances per symbolic template with distinct random seeds, yielding 58 topics $\times$ 5 templates $\times$ 10 instances = 2,900 test cases. All models are evaluated under a unified decoding configuration: temperature = 0.7, top-p = 0.95, and a maximum token limit of 4,096 unless these parameters are unavailable for a given model. We use a zero-shot setup with a standardized reasoning prompt:
@@
We use ChainEval as the primary evaluation metric, as it jointly measures final-answer correctness and alignment of intermediate reasoning steps. Model outputs are post-processed with regular expressions to extract the ordered list of reasoning steps, accommodating common variations such as “Step x:”, “Step x”, or “stepx”. For comparison, we also report frequently used reference-based metrics (Xie et al., 2023c) to assess surface-level quality, including ROUGE-2 and ROUGE-L (Lin, 2004) and BERTScore (Zhang et al., 2020). We evaluate most models using the full test set, and the main results are presented in Table 1. However, due to the high cost of Grok4 Heavy (approximately \$0.5 per sample), we randomly sample 200 test cases and compare its performance with the top-performing models, as shown in Table 2.
5.4 Results and Analysis
@@ 5.4.1 Overall Model Performance
Model Size ChainEval $\uparrow$ ROUGE R2 $\uparrow$ ROUGE RL $\uparrow$ BERTScore $\uparrow$
@@
GPT-5 N/A 57.077.28 28.8412.30 42.7712.91 88.772.39
GPT-4.1 N/A 56.926.12 19.388.91 30.1210.57 86.042.09
GPT-5-mini N/A 57.387.12 26.4811.99 39.7412.83 88.182.35
GPT-4.1-mini N/A 57.245.82 18.678.86 29.0510.51 86.052.09
Claude Sonnet 4.5 N/A 58.226.43 19.697.77 29.378.81 86.072.00
Claude Sonnet 4 N/A 58.186.60 19.877.82 29.508.59 86.381.83
Claude Sonnet 3.7 N/A 57.896.22 19.497.77 29.368.56 86.381.82
Gemini 2.5 Pro N/A 58.656.87 17.617.07 27.368.28 85.941.88
Gemini 2.5 Flash N/A 58.016.99 18.988.05 29.229.32 86.342.02
DeepSeek v3.2 N/A 56.716.25 21.7310.32 32.8511.31 86.662.15
DeepSeek v3.1 N/A 56.766.40 21.7210.29 32.8711.24 86.682.14
DeepSeek R1 N/A 53.757.90 8.677.21 12.939.72 84.391.79
Grok 4 Fast N/A 56.7315.59 21.3311.09 32.2513.30 86.833.06
Finance Specific LLMs
Fin-o1 8B 39.3412.10 3.471.55 6.352.32 83.551.50
Fin-R1 7B 57.345.62 5.702.44 9.223.33 84.301.34
DianJin-R1 7B 53.727.61 6.282.95 10.794.19 83.121.32
Finance-LlaMA 8B 42.819.33 9.394.69 16.195.84 83.482.09
Finance-Qwen 7B 34.2210.69 9.504.26 16.465.44 83.351.70
Math Enhanced LLMs
WizardMath 7B 21.7515.45 11.666.57 20.727.83 84.782.36
MetaMath 7B 6.099.24 11.457.36 21.089.24 84.862.99
Mathstral 7B 56.405.96 16.797.82 26.979.34 86.132.18
Qwen2.5-Math 7B 50.3212.64 11.745.87 20.567.61 83.451.85
General Purpose Open LLMs
Llama 3.1 Instruct 8B 55.884.95 4.612.28 8.093.02 83.351.36
Qwen 2.5 Instruct 7B 57.005.35 9.204.51 15.265.85 84.221.78
Qwen 3 8B 45.9910.84 4.051.69 6.612.14 83.581.24
@@
Model ChainEval $\uparrow$ ROUGE R2 $\uparrow$
Grok 4 Heavy 61.99 23.87
GPT-5 60.60 28.37
Gemini 2.5 Pro 58.47 18.85
Fin-R1 57.67 5.43
Mathstral 57.15 18.99
Table 2: Performance comparison over 200 random samples using ChainEval and ROUGE R2, considering they are the closest metrics to human evaluation according to Table 5.
As shown in Table 1, frontier proprietary models such as GPT-5, Claude 4.5, and Gemini 2.5 Pro dominate across all metrics, achieving the highest ChainEval, ROUGE, and BERTScore values. Their advantage reflects stronger symbolic reasoning and alignment quality, driven by larger model capacity and more comprehensive instruction tuning. Yet the gap with open models is narrowing: Fin-R1 and Mathstral reach ChainEval scores of 57.34 and 56.40, approaching frontier performance despite being only 7B in scale. This highlights that strategic fine-tuning, rather than sheer size, can substantially improve multi-step reasoning fidelity. However, the benefits of fine-tuning hinge on adaptation scope. Models trained on broad mathematical corpora (e.g., Mathstral) generalize better across reasoning styles, while those tuned narrowly for finance (e.g., Finance-LLaMA, Finance-Qwen) gain domain accuracy at the cost of generality. This trade-off suggests that future financial LLMs should balance symbolic supervision and domain grounding to achieve both precision and transferability. Table 2 further supports these trends: on 200 sampled cases, Grok 4 Heavy achieves the highest ChainEval (61.99), narrowly surpassing GPT-5 (60.60), though GPT-5 yields the best ROUGE R2 (28.37), indicating closer alignment with human reasoning traces. Meanwhile, Fin-R1 (57.67) and Mathstral (57.15) remain competitive, confirming that fine-tuned and math-enhanced open models can approximate frontier-level reasoning when guided by high-quality symbolic supervision.
5.4.2 Performance Over Domains
@@
Figure 3: Domain-level performance across financial domains. Radar plot comparing the best-performing model from each category in Table 1, evaluated across twelve financial domains using ChainEval scores.
Figure 3 illustrates domain-level ChainEval performance across twelve financial categories, comparing the best proprietary and fine-tuned open models. The frontier system GPT-5 maintains consistently strong and balanced performance across nearly all domains, reflecting broad reasoning coverage and stable generalization. Gemini 2.5 Pro exhibits a similar trend but shows slight advantages in corporate finance, fintech, and financial markets, likely benefiting from broader exposure to business-oriented content during pretraining. Together, these results confirm that large-scale proprietary LLMs continue to offer the most uniform and reliable performance across diverse financial contexts.
In contrast, fine-tuned open models display domain-selective strengths that align with their adaptation objectives. Fin-R1 achieves competitive or superior accuracy in financial reporting, sustainable finance, and risk management, outperforming frontier models in several specialized areas. Mathstral performs strongly in quantitative domains such as financial ratios and investment analysis but trails in text-heavy categories like policy or corporate finance. These asymmetric patterns indicate that fine-tuning effectively amplifies reasoning in domain-relevant tasks but introduces variability across unrelated ones. Overall, fine-tuned open models demonstrate meaningful progress toward domain robustness, yet frontier proprietary systems remain more balanced and reliable across the entire financial spectrum.
5.4.3 Performance Over Difficulty Levels
Figure 4: ChainEval score across difficulty levels.
To assess model robustness under increasing complexity of reasoning, we group the results into three predefined tiers (Basic, Intermediate, and Advanced (§ 3). Each difficulty level reflects an increase in reasoning and computational depth. Figure 4 presents model performance across three difficulty tiers measured by ChainEval scores. Across all tiers, frontier proprietary models demonstrate the highest overall correctness, with GPT-5 and Gemini 2.5 Pro showing strong resilience as task complexity increases. Their relatively small performance drop between Basic and Advanced levels highlights more stable reasoning and abstraction capabilities, likely reflecting the benefits of extensive instruction alignment and large, diverse training corpora. These trends reaffirm that frontier models maintain superior robustness and generalization across varying levels of reasoning difficulty.
In contrast, open and fine-tuned models exhibit sharper performance declines as difficulty increases but still show meaningful gains relative to their base architectures. Mathstral achieves competitive scores on Basic and Intermediate tasks, demonstrating that mathematical fine-tuning enhances structured reasoning, while Fin-R1 performs comparatively well on simpler finance-related queries but degrades more on Advanced ones. This suggests that fine-tuning improves domain reasoning and factual accuracy but offers limited transfer to high-complexity, multi-step tasks. Overall, the results indicate that while fine-tuned open models narrow the gap at lower difficulty levels, frontier proprietary systems retain a clear advantage in handling complex, compositional reasoning.
6 Conclusions and Future Work
We introduce FinChain, a symbolic benchmark for verifiable chain-of-thought financial reasoning. Built from a fine-grained taxonomy spanning 58 topics across 12 domains and three difficulty levels, it enables targeted evaluation across reasoning complexity and domain-specific challenges. To assess performance, we propose ChainEval, a dynamic alignment metric evaluating both final-answer correctness and intermediate-step consistency. Experiments show that while frontier LLMs achieve the highest accuracy, even the strongest systems struggle with complex symbolic reasoning. Fine-tuned open-source models such as Fin-R1 and Mathstral reduce but do not close this gap, underscoring the challenge of aligning fluency with verifiable reasoning. Future work will expand FinChain to multilingual and region-specific settings and explore how explicit reasoning traces can improve the trustworthiness of model-generated financial analyses, particularly in long-form question answering over real documents. This direction aims to bridge symbolic reasoning and factual verification (Xie et al., 2025), advancing more interpretable and reliable financial AI systems.

### v3

@@ 4.3 Final Answer Correctness
Beyond step-level reasoning alignment, we also assess the correctness of the final predicted outcome. Let $s^{*}_{m}$ and $\hat{s}_{n}$ denote the last steps of the gold and the predicted solutions, respectively. We define the Final Answer Correctness (FAC) metric as
@@
using the same tolerance $\epsilon=0.05$ as before. FAC measures whether the model’s final computation aligns with the correct end result, complementing the DTW-based metric that evaluates reasoning faithfulness throughout the entire solution sequence. Therefore, we have
$\textsc{ChainEval}=(1-\alpha)\cdot\text{DTWGate}+\alpha\cdot\text{FAC},$ (8)
which accounts for both reasoning correctness and final answer’s correctness. We set $\alpha=0.1$, selected via grid search on a subset to maximize the correlation with human evaluations, as explained in Appendix D. We further verify that the final measure best reflects true reasoning quality by comparing it with human evaluations, as described in § 5.2.
@@ 5.1 Evaluated Models
We evaluate a total of 26 LLMs, grouped into four categories according to their capability and relevance to financial reasoning. (1) Frontier proprietary models, including GPT-{5, 4.1, 5-mini, 4.1-mini} (OpenAI, 2025a, b), Claude Sonnet {4.5, 4, 3.7} (Anthropic, 2025b, c, a), Gemini 2.5 {Pro, Flash} (Comanici et al., 2025), DeepSeek {V3.2, V3.1, R1} (Liu et al., 2024; Guo et al., 2025), and Grok 4 {Heavy, Fast} (xAI, 2025). (2) Finance-specific models, including Fino1 (Qian et al., 2025), FinR1 (Liu et al., 2025), DianJin-R1 (Zhu et al., 2025), and WiroAI Finance-{LLaMA, Qwen} (Abdullah Bezir, 2025). (3) Math-enhanced models, including WizardMath (Luo et al., 2023), MetaMath (Yu et al., 2023), Mathstral (Mistral, 2024), and Qwen2.5-Math (Yang et al., 2024). (4) General-purpose open-weight models, including LLaMA 3.1 (Grattafiori et al., 2024) and Qwen {2.5, 3} (Qwen, 2024, 2025). Detailed configurations and model sources are described in Appendix C.
@@ 5.2 ChainEval Validation
Before conducting large-scale experiments, we validated the proposed ChainEval through a controlled expert evaluation. We randomly sampled 20 instantiated questions from the dataset and generated answers using five models of different capacities and training paradigms, namely GPT-5, GPT-4.1 mini, MetaMath, Fino1, and LLaMA 3.1, to ensure a clear range of reasoning quality. This process produced 100 model-generated responses, which were then randomized and anonymized for human assessment. Financial experts independently evaluated each response with respect to the corresponding question and gold-standard reasoning trace. They rated each output along two dimensions, Reasoning Process Quality and Final Answer Accuracy, on a five-point scale as described in Appendix B.4. We observed a strong association between these dimensions, with Spearman’s $\rho$ exceeding 0.94, showing that coherent reasoning typically leads to accurate final answers. Consequently, subsequent analysis below will focus on Reasoning Process Quality as the primary evaluation dimension. We further computed several variants of ChainEval and compared their correlations with expert process scores against reference-based evaluation measures such as ROUGE-2 and ROUGE-L (Lin, 2004). Among all variants, the DTWNormGate formulation with inclusion of final answer’s correctness showed the highest correlation with human judgments, capturing both semantic and numeric consistency across reasoning steps. Therefore, we adopted it as the primary evaluation measure in this study, with detailed ablation results reported in Appendix D.
@@ 5.3 Experimental Setup
We instantiate the benchmark by sampling 10 instances per symbolic template with distinct random seeds, yielding 58 topics $\times$ 5 templates $\times$ 10 instances = 2,900 test cases. We evaluated all models under a unified decoding configuration: temperature = 0.7, top-p = 0.95, and a maximum token limit of 4,096 unless these parameters are unavailable for a given model. We use a zero-shot setup with a standardized reasoning prompt:
@@
We use ChainEval as the primary evaluation measure, as it jointly measures final-answer correctness and alignment of the intermediate reasoning steps. We post-processed model outputs with regular expressions to extract the ordered list of reasoning steps, accommodating common variations such as “Step x:”, “Step x”, or “stepx”. For comparison, we also report results using frequently used reference-based measures (Xie et al., 2023c) to assess surface-level quality, including ROUGE-2 and ROUGE-L and BERTScore (Zhang et al., 2020).
5.4 Results
@@ 5.4.1 Overall Model Performance
Model Size ChainEval $\uparrow$ FAC $\uparrow$ ROUGE R2 $\uparrow$ ROUGE RL $\uparrow$ BERTScore $\uparrow$
@@
GPT-5 N/A 66.5710.64 82.0332.40 28.8412.30 42.7712.91 88.772.39
GPT-4.1 N/A 65.349.36 84.6629.26 19.388.91 30.1210.57 86.042.09
GPT-5-mini N/A 67.1711.06 80.2832.34 26.4811.99 39.7412.83 88.182.35
GPT-4.1-mini N/A 65.068.88 84.5930.23 18.678.86 29.0510.51 86.052.09
Claude Sonnet 4.5 N/A 66.339.44 83.3431.79 19.697.77 29.378.81 86.072.00
Claude Sonnet 4 N/A 66.209.66 82.6231.96 19.877.82 29.508.59 86.381.83
Claude Sonnet 3.7 N/A 65.519.30 83.1431.00 19.497.77 29.368.56 86.381.82
Gemini 2.5 Pro N/A 66.049.71 84.3430.99 17.617.07 27.368.28 85.941.88
Gemini 2.5 Flash N/A 65.9610.10 83.9030.61 18.988.05 29.229.32 86.342.02
DeepSeek v3.2 N/A 65.239.63 84.1731.23 21.7310.32 32.8511.31 86.662.15
DeepSeek v3.1 N/A 65.299.62 84.3431.02 21.7210.29 32.8711.24 86.682.14
DeepSeek R1 N/A 51.2211.04 28.9730.29 8.677.21 12.939.72 84.391.79
Grok 4 Fast N/A 60.6916.26 66.5442.54 21.3311.09 32.2513.30 86.833.06
Finance-Specific LLMs
Fin-o1 8B 41.5012.49 52.7927.89 3.471.55 6.352.32 83.551.50
Fin-R1 7B 58.147.32 52.7628.04 5.702.44 9.223.33 84.301.34
DianJin-R1 7B 51.958.98 37.6923.73 6.282.95 10.794.19 83.121.32
Finance-LLaMA 8B 41.3510.49 25.2125.64 9.394.69 16.195.84 83.482.09
Finance-Qwen 7B 34.5711.01 31.6225.62 9.504.26 16.465.44 83.351.70
Math-Enhanced LLMs
WizardMath 7B 24.3315.00 41.2835.69 11.666.57 20.727.83 84.782.36
MetaMath 7B 7.939.43 23.9728.76 11.457.36 21.089.24 84.862.99
Mathstral 7B 59.8710.02 54.0336.61 16.797.82 26.979.34 86.132.18
Qwen2.5-Math 7B 55.3514.98 62.6234.56 11.745.87 20.567.61 83.451.85
General-Purpose Open LLMs
LLaMA 3.1 Instruct 8B 53.996.02 32.7227.46 4.612.28 8.093.02 83.351.36
Qwen 2.5 Instruct 7B 60.357.47 65.4132.53 9.204.51 15.265.85 84.221.78
Qwen 3 8B 43.3211.81 32.2828.58 4.051.69 6.612.14 83.581.24
@@
As shown in Table 2, frontier proprietary models achieve the strongest overall performance under ChainEval, indicating higher alignment quality in step-level symbolic reasoning. The performance differences between frontier models are not strictly monotonic. For example, GPT-5-mini slightly outperforms GPT-5, despite comparable or lower scores on surface-level metrics. Among open-source systems, Fin-R1 and Mathstral achieve ChainEval scores of 58.14 and 59.87 respectively, approaching frontier-level alignment despite their substantially smaller scale. These results suggest that targeted fine-tuning and symbolic supervision can substantially improve multi-step reasoning fidelity beyond what model size alone provides. However, the effectiveness of fine-tuning varies with adaptation scope: models trained on broad mathematical corpora (e.g., Mathstral) exhibit stronger generalization across reasoning styles, whereas models tuned narrowly for finance (e.g., Finance-LLaMA and Finance-Qwen) display more variable alignment quality. We also evaluate Grok 4 Heavy, which was among the strongest available systems at the time of evaluation. Due to its high inference cost (approximately \$0.5 per sample), we only assessed this system on a randomly sampled subset rather than the full benchmark, with results reported in Appendix E.1. Overall, these results show that while frontier models lead in verifiable symbolic reasoning, model scale alone is insufficient, and structured supervision plays a critical role in achieving high step-level alignment.
5.4.2 Performance Across Domains
@@
Figure 3: Domain-level performance across financial domains. Radar plot showing ChainEval scores across twelve financial domains for four representative models: GPT-5, Gemini 2.5 Pro, Fin-R1, and Mathstral.
Figure 3 shows the domain-level ChainEval performance across twelve financial categories for four representative models: GPT-5, Gemini 2.5 Pro, Fin-R1, and Mathstral. Among them, GPT-5 demonstrates consistently strong performance across most domains, forming a relatively stable upper envelope with limited variation. Gemini 2.5 Pro follows a similar overall trend, achieving competitive scores across domains while exhibiting moderate domain-level differences. The two open-weight models show more heterogeneous domain-specific patterns. Fin-R1 achieves relatively higher scores in domains such as Financial Reporting, Sustainable Finance, and Risk Management, while exhibiting lower performance in several quantitatively intensive or structurally complex domains. In contrast, Mathstral performs competitively in quantitatively orientated domains, including Financial Ratios and Investment Analysis, but shows weaker performance in domains with heavier textual or regulatory components. In general, domain-level performance is not uniform across models and relative rankings vary by financial category. These observations highlight systematic variation across domains and motivate further diagnostic analysis in subsequent sections. Additional domain-level results for the remaining models are provided in Appendix E.2. Overall, these results indicate that symbolic financial reasoning ability is highly domain-dependent, and that strong aggregate performance does not guarantee robust reasoning across diverse financial categories.
5.4.3 Performance Across Difficulty Levels
Figure 4: Final answer correctness (FAC) across difficulty levels. FAC for representative models on Basic, Intermediate, and Advanced FinChain instances.
To assess the model robustness under increasing reasoning complexity, we group FinChain instances into three predefined difficulty tiers: Basic, Intermediate, and Advanced (§ 3). Each tier corresponds to an increase in the number of required reasoning steps and the depth of symbolic and numerical operations. Figure 4 reports model performance across difficulty levels using final answer correctness (FAC), isolating end-task success from partial step-level alignment captured by ChainEval. Across all tiers, frontier proprietary models achieve the highest correctness, with GPT-5 and Gemini 2.5 Pro maintaining relatively strong performance as task difficulty increases. Nevertheless, even these models exhibit a clear degradation on advanced instances, highlighting the challenge of solving complex, multi-step financial reasoning problems to completion. In contrast, open-weight and fine-tuned models show substantially steeper declines as difficulty increases. Mathstral performs competitively on Basic and Intermediate tasks, suggesting that mathematical fine-tuning improves structured numerical reasoning, but its performance drops markedly on Advanced instances. Fin-R1 displays similar behavior, achieving reasonable accuracy on simpler finance-oriented queries while degrading more sharply on tasks requiring longer reasoning chains. Overall, FAC reveals a pronounced gap between frontier and non-frontier models on advanced symbolic reasoning tasks. Taken together with the more gradual degradation observed under ChainEval, these results suggest that completing long, multi-step reasoning chains remains a key bottleneck. We report the corresponding difficulty breakdown under ChainEval in Appendix E.3.
5.5 Error Analysis
Figure 5: Error type distribution on a randomly sampled subset of 200 questions.
We conduct a targeted error analysis on a randomly sampled subset of 200 questions from FinChain, using the same question set for GPT-5-mini and Finance-Qwen. These two models serve as contrasting points on the performance spectrum under an identical evaluation setup, with GPT-5-mini as a strong frontier-aligned reference and Finance-Qwen as a finance-tuned model with lower overall accuracy. This shared subset enables a direct comparison of error profiles without confounding effects from domain coverage or evaluation scale. We manually inspect model outputs and assign them to a small set of coarse-grained error categories capturing common reasoning failures in financial problem solving. Our taxonomy includes Computational errors (incorrect numerical execution), Conceptual errors (incorrect application of financial concepts), Sign/Direction errors, Unit/Precision mismatches, Parsing/Format issues that prevent reliable step alignment, and Hallucination, which includes content unsupported by the input. Outputs that are fully correct within tolerance are labeled as No Error. A detailed description of each category is provided in Appendix F.1. During expert review, a small number of borderline cases initially categorized as computational errors were revised to No Error after manual inspection, primarily when numerical deviations were deemed immaterial. All reported statistics reflect these expert-validated labels. Figure 5 summarizes the resulting error distributions. GPT-5-mini produces correct solutions for most cases, with remaining failures concentrated in numerical computation and conceptual reasoning. Other error types are rare, and no hallucinated or unparsable outputs are observed on this subset. In contrast, Finance-Qwen exhibits errors on most evaluated cases, with failures spread across multiple categories. Conceptual and computational errors dominate, while a substantial fraction involves formatting issues or hallucinated information. Compared to GPT-5-mini, Finance-Qwen errors are less concentrated in a single mode and reflect more diverse failure behaviors. Examples of representative errors are provided in Appendix F.2. Overall, this analysis shows that symbolic financial reasoning errors are heterogeneous and model-dependent, with weaker models exhibiting more diverse and compounding failure modes, helping explain the aggregate performance gaps.
6 Conclusion and Future Work
We introduced FinChain, a symbolic benchmark for verifiable Chain-of-Thought financial reasoning, spanning 58 topics across 12 domains and three difficulty levels. To support step-level evaluation, we proposed ChainEval, which jointly assesses intermediate reasoning consistency and final-answer correctness. Our results showed that while frontier LLMs perform best overall, even the strongest models struggle with complex symbolic reasoning, and fine-tuned open-source systems narrow down, but do not close this gap.
In future work, we plan to extend FinChain to multilingual and region-specific settings and to incorporate problems grounded in real-world financial documents. This direction aims to bridge symbolic reasoning and factual verification (Xie et al., 2025), advancing more interpretable and reliable financial AI systems.